Light of wavelength 519 nm passes through two slits. In the interference pattern on a screen 4.6 m away, adjacent bright fringes are separated by 5.2 mm in the general vicinity of the center of the pattern. What is the separation of the two slits?

Draw the slits • Draw the screen a distance L from the slits • Draw the paths from each slit • Mark the bright locations on the screen.

Start with the double slit bright fringe equation • Substitute in the known values • Solve for the desired unknown • Check units and assess if your answer is reasonable
Answer #1

Wavelength of light, λ = 519 x 10-9 m

Distance of the screen from the slits, D = 4.6 m

Distance between two adjacent bright fringes (fringe width), β = 5.2 mm = 5.2 x 10-3 m

Let d be the separation between two slits.

Fringe width in interference pattern, β = λD/d

the separation between two slits, d = λD/β = 519 x 10-9 x 4.6/5.2 x 10-3 = 4.50 x 10-4 m

the separation between two slits, d = 4.50 x 10-4 m